Consider the integral "\displaystyle{\int \frac{2}{x^2-2x}dx}=\displaystyle{\ln{\frac{x-a}{x}}+C}"

 Which of the following values represent the constant a ?

 

  1. 1
  2. 7
  3. 4
  4. 2




1
Expert's answer
2021-09-06T16:47:02-0400

Solve the integral by partial fractions


"\dfrac{2}{x^2-2x}=\dfrac{2}{x(x-2)}=\dfrac{A}{x}+\dfrac{B}{x-2}"

The sum is added up:


"\dfrac{2}{x^2-2x}=\dfrac{A(x-2)+Bx}{x(x-2)}"

Since the two fractions have the same denominator, the numerators must be equal:


"2=A(x-2)+Bx"

To calculate the values of "A" , "B" , we give "x" the values that cancel out the denominator


"x=0\Rightarrow 2=-2A\Rightarrow A=-1"

"x=2\Rightarrow 2=2B\Rightarrow B=1"

Integrals of simple fractions are calculated:


"\int \dfrac{2}{x^2-2x}\, dx=-\int \dfrac1x\, dx+\int \dfrac{1}{x-2}\, dx"

"=-\ln |x|+\ln |x-2|+C=\ln \left| \dfrac{x-2}{x}\right|+C"

Comparing with the expression of the statement we see that "\boxed{a=2}."